OpenStudy (anonymous):

i have this question i need help with. Write an equation that represents the line that passes through the points (5,4) and (-5,0)?

OpenStudy (anonymous):

first use the slope formula (y2-y1)/(x2-x1), or (0-4)/(-5-5)=-4/-10=2/5

OpenStudy (anonymous):

after this you can use the slope-point formula y-y1=m(x-x1) or y-4=2/5(x-5)